(talking about the apollo spacecraft, ignore air resistance) if the 28900kg Apollo 9 CSM sprays 2kg of urine out with a velocity of 20 m/s

a. what velocity would this give the spacecraft

b. over a period of 24 hours, by what distance would this change Apollo’s path?


1
Expert's answer
2021-01-28T20:10:53-0500

According to the momentum conservation law, the amount of momentum carried out by urine:


"p_u = 2kg\times 20m/s = 40kg\cdot m/s"

is equal to the change of Apollo's momentum:


"m_a\Delta v_a = p_u"

where "m_a = 28900kg" is the mass of Apollo (neglecting the change of mass due to 2kg urine left), and "\Delta v_a" is the change of its velocity:


"\Delta v_a = \dfrac{p_u}{m_a}\\
\Delta v_a = \dfrac{40kg\cdot m/s}{28900kg} \approx 0.00138m/s"

Over the period "t = 24h = 86400s" this would change Apollo’s path by the following distance:


"\Delta d = \Delta v_at\\
\Delta d = 0.0014m/s\cdot 86400s \approx 119m"

Answer. a) 0.00138m/s, b) 119m.